OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4.

OpenStudy (anonymous):

Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

\[Point \space Slope \space Form: y - y_{1} = m(x - x_{1})\]\[Slope-Intercept \space Form: y = mx + b\]\[Standard \space Form: Ax + By = C\]

OpenStudy (anonymous):

thnks that helps alot !

OpenStudy (anonymous):

np :) If you want me to check your work, just post your answers here and I'll check.